OpenStudy (anonymous):

two ships leave a port at 9am. one travels at a bearing of N 53 degrees W at 12 miles per hour, and the other travels at a bearing of S 67 degrees W at 16 miles per hour. Approximately how far apart there are at noon that day.

OpenStudy (zpupster):



OpenStudy (zpupster):

step 1 is find sides of the triangle 12 m/h and 16 m/h for how three hrs a = ? b=? what is the angle between them?? remember we have the vertical axis at 180 so 180 - 53-67=?? the use the law of cosine c(or side we want) c^2 = a^2xb^2-2abcos( of angle between)